# Learning Performance Maximizing Ensembles with Explainability Guarantees

### v2

@@ Appendix A Appendix
Part 1: We show that $b^{*}_{q}\in B^{*}_{q}\implies b^{*}_{q}\in A^{*}_{q}$ for given $q\in\mathcal{Q}$. Assume for contradiction that $b^{*}_{q}\notin A^{*}_{q}$. Then $\exists(z_{i}^{\circ},z_{j}^{\circ})\in\mathcal{Z}\times\mathcal{Z}$ s.t. $[s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q}(z_{i}^{\circ})+s^{\prime}_{b}(z_{i}^{\circ})(1-b^{*}_{q}(z_{i}^{\circ}))]+[s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q}(z_{j}^{\circ})+s^{\prime}_{b}(z_{j}^{\circ})(1-b^{*}_{q}(z_{j}^{\circ}))]<[s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q}(z_{j}^{\circ})+s^{\prime}_{b}(z_{i}^{\circ})(1-b^{*}_{q}(z_{j}^{\circ}))]+[s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q}(z_{i}^{\circ})+s^{\prime}_{b}(z_{j}^{\circ})(1-b^{*}_{q}(z_{i}^{\circ}))]$ i.e. there exists a pair of observations for which swapping allocation decisions increases the objective function.
Case 1 (L.3.1.1): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{02,02}$. But, $\forall(z_{i},z_{j})\in Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{02,02}$, we have that $s^{\prime}_{g}(z_{i})b^{*}_{q}(z_{i})+s^{\prime}_{b}(z_{i})(1-b^{*}_{q}(z_{i}))]+[s^{\prime}_{g}(z_{j})b^{*}_{q}(z_{j})+s^{\prime}_{b}(z_{j})(1-b^{*}_{q}(z_{j}))]=[s^{\prime}_{g}(z_{i})b^{*}_{q}(z_{j})+s^{\prime}_{b}(z_{i})(1-b^{*}_{q}(z_{j}))]+[s^{\prime}_{g}(z_{j})b^{*}_{q}(z_{i})+s^{\prime}_{b}(z_{j})(1-b^{*}_{q}(z_{i}))]$ which contradicts the definition of $(z_{i}^{\circ},z_{j}^{\circ})$, thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{02,02}$.
Case 2 (L.3.1.2): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,02}$. Then, $t^{\prime}(b^{*}_{q},z_{i}^{\circ})=b^{*}_{q}(z_{i}^{\circ}),t^{\prime}(b^{*}_{q},z_{j}^{\circ})=c=\begin{cases}1,z_{j}^{\circ}\in\mathcal{Z}^{\prime}_{2}\\ 0,z_{j}^{\circ}\in\mathcal{Z}^{\prime}_{0}\end{cases}$. Thus, $t^{\prime}(b^{*}_{q},z_{i}^{\circ})+t^{\prime}(b^{*}_{q},z_{j}^{\circ})=b^{*}_{q}(z_{i}^{\circ})+c$ which by assumption on $z_{i}^{\circ},z_{j}^{\circ}\implies b^{*}_{q}(z_{i}^{\circ})=0<b^{*}_{q}(z_{j}^{\circ})=1\implies C_{g,02}(b^{*}_{q})\neq\emptyset$ which is a contradiction with $b^{*}_{q}\in B^{*}_{q}$, so $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{g,02}$.
Case 3 (L.3.1.3): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{02,b}$. Using the same argument as in Case 2 (L.3.1.2), contradiction and thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{02,b}$ follows.
Case 4 (L.3.1.4): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,b}$. Using the same argument as in Case 2 (L.3.1.2), contradiction and thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{g,b}$ follows.
Thus we have shown that $\forall q$, no such $z_{i}^{\circ},z_{j}^{\circ}$ exist, and so $b^{*}_{q}\in A^{*}_{q}$.
Part 2: We show $a^{*}_{q}\in A^{*}_{q}\implies a^{*}_{q}\in B^{*}_{q}$ for given $q\in\mathcal{Q}$. Assume for contradiction that $a^{*}_{q}\notin B^{*}_{q}$. Then $\exists z_{i}^{\circ},z_{j}^{\circ}$ s.t. $(z^{\circ}_{i},z^{\circ}_{j})\in C_{g,02}\cup C_{02,b}\cup C_{g,b}$.
Case 1 (L.3.2.1): $(z_{i}^{\circ},z_{j}^{\circ})\in C_{g,02}$. Then, $t^{\prime}(a^{*}_{q},z_{i}^{\circ})=a_{q}(z_{i}^{\circ})$ and $t^{\prime}(a^{*}_{q},z_{j}^{\circ})=c=\begin{cases}1,z_{j}^{\circ}\in\mathcal{Z}^{\prime}_{2}\\ 0,z_{j}^{\circ}\in\mathcal{Z}^{\prime}_{0}\end{cases}$ and $a^{*}_{q}(z_{i}^{\circ})=0<1=a^{*}_{q}(z_{j}^{\circ})$. But then, $t^{\prime}(a^{*}_{q},z_{i}^{\circ})+t^{\prime}(a^{*}_{q},z_{j}^{\circ})=c<1+c=[s^{\prime}_{g}(z_{i}^{\circ})a^{*}_{q}(z_{j}^{\circ})+s^{\prime}(z_{i}^{\circ})(1-a^{*}_{q}(z_{j}^{\circ}))]+[s^{\prime}_{g}(z_{j}^{\circ})a^{*}_{q}(z_{i}^{\circ})+s^{\prime}(z_{j}^{\circ})(1-a^{*}_{q}(z_{i}^{\circ}))]$ i.e. swapping the allocation decisions of $z_{i}$ and $z_{j}$ increases the objective function $\implies a^{*}_{q}\notin A^{*}_{q}$ which is a contradiction. Thus, $(z_{i}^{\circ},z_{j}^{\circ})\notin C_{g,02}$.
Case 2 (L.3.2.2): $(z_{i}^{\circ},z_{j}^{\circ})\in C_{02,b}$. Using the same argument as in Case 1 (L.3.2.1), $z_{i}^{\circ},z_{j}^{\circ}\notin C_{02,b}$ follows.
Case 3 (L.3.2.3): $(z_{i}^{\circ},z_{j}^{\circ})\in C_{g,b}$. Using the same argument as in Case 1 (L.3.2.1), $z_{i}^{\circ},z_{j}^{\circ}\notin C_{g,b}$ follows.
Thus, $(z^{\circ}_{i},z^{\circ}_{j})\notin C_{g,02}\cup C_{02,b}\cup C_{g,b}\implies a^{*}_{q}\in B^{*}_{q}$
Having shown both directions of implication, we conclude $A^{*}_{q}=B^{*}_{q}$
@@
(Maximal Sufficient Performance) $\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q}$ where $A^{*}_{q}=\{a^{*}_{q}\colon a^{*}_{q}=\arg\max_{a_{q}\in A_{q}}\bar{t}^{\prime}(a_{q})\}$ and $\bar{t}^{\prime}(a_{q})=\frac{1}{n}\sum_{i=1}^{n}t^{\prime}(a_{q},z_{i})=\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a_{q}(z_{i})+s^{\prime}_{b}(z_{i})(1-a_{q}(z_{i}))$
@@
From Lemma 3, it is enough to show that $a^{\prime}_{q}\in B^{*}_{q}$. First, from Lemma 2, it follows that $a^{\prime}_{q}\in A_{q}$. Next, we show that $C_{g,02}(a^{\prime}_{q})=C_{02,b}(a^{\prime}_{q})=C_{g,b}(a^{\prime}_{q})=\emptyset$.
Case 1 (P.1.1): $(z^{\circ}_{i},z^{\circ}_{j})\in Z^{\prime}_{g,02}$. From Lemma 1, $\tilde{r}^{\prime}(z^{\circ}_{i})>\tilde{r}^{\prime}(z^{\circ}_{j})\iff\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z^{\circ}_{i}))>\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z^{\circ}_{j}))\implies\mathbb{I}\{\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z^{\circ}_{i}))>n_{1-q}\}\geq\mathbb{I}\{\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z^{\circ}_{j}))>n_{1-q}\}\iff a^{\prime}_{q}(z^{\circ}_{i})\geq a^{\prime}_{q}(z^{\circ}_{j})$ and thus $(z^{\circ}_{i},z^{\circ}_{j})\notin C_{g,02}(a^{\prime}_{q})$.
Case 2 (P.1.2): $(z^{\circ}_{i},z^{\circ}_{j})\in Z^{\prime}_{02,b}$. Using the same argument as in Case 1 (P.1.1), $(z^{\circ}_{i},z^{\circ}_{j})\notin C_{02,b}(a^{\prime}_{q})$ follows.
Case 3 (P.1.3): $(z^{\circ}_{i},z^{\circ}_{j})\in Z^{\prime}_{g,b}$. Using the same argument as in Case 1 (P.1.1), $(z^{\circ}_{i},z^{\circ}_{j})\notin C_{g,b}(a^{\prime}_{q})$ follows.
Case 4 (P.1.4): $(z^{\circ}_{i},z^{\circ}_{j})\in(\mathcal{Z}\times\mathcal{Z})-(Z^{\prime}_{g,02}\cup Z^{\prime}_{02,b}\cup Z^{\prime}_{g,b})\implies(z^{\circ}_{i},z^{\circ}_{j})\notin C_{g,02}(a^{\prime}_{q})\cup C_{02,b}(a^{\prime}_{q})\cup C_{g,b}(a^{\prime}_{q})$ by definition.
We have shown that $a^{\prime}_{q}\in A_{q}$ and $C_{g,02}(a^{\prime}_{q})=C_{02,b}(a^{\prime}_{q})=C_{g,b}(a^{\prime}_{q})=\emptyset$, thus $a^{\prime}_{q}\in B^{*}_{q}=A^{*}_{q}$. ∎
Lemma 4.
(Set Equivalence 2) $A^{*}_{q|g}=B^{*}_{q|g}$ where $q\in\mathcal{Q},A^{*}_{q|g}=\{a^{*}_{q|g}\colon a^{*}_{q|g}=\arg\max_{a^{*}_{q}\in A^{*}_{q}}\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a^{*}_{q}(z_{i})\}$ and $B^{*}_{q|g}=\{b^{*}_{q|g}:b^{*}_{q|g}\in A^{*}_{q},C_{2,0}(b^{*}_{q|g})=\emptyset\}$
@@
Part 1: First we show that $b^{*}_{q|g}\in B^{*}_{q|g}\implies b^{*}_{q|g}\in A^{*}_{q|g}$. Assume for contradiction that $b^{*}_{q|g}\notin A^{*}_{q|g}$, then $\exists(z_{i}^{\circ},z_{j}^{\circ})\in\mathcal{Z}\times\mathcal{Z}$ s.t. $\alpha=s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q|g}(z_{i}^{\circ})+s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q|g}(z_{j}^{\circ})<s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q|g}(z_{j}^{\circ})+s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q|g}(z_{i}^{\circ})=\beta$ i.e. there exists a pair of observations for which swapping allocation decisions increases the $g$-specific objective function. Next, $b^{*}_{q|g}\in B^{*}_{q|g}\implies b^{*}_{q|g}\in A^{*}_{q}\implies\forall(z_{i},z_{j})\in\mathcal{Z}\times\mathcal{Z}$, and in particular for $(z_{i}^{\circ},z_{j}^{\circ})$, $[s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q|g}(z_{i}^{\circ})+s^{\prime}_{b}(z_{i}^{\circ})(1-b^{*}_{q|g}(z_{i}^{\circ}))]+[s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q|g}(z_{j}^{\circ})+s^{\prime}_{b}(z_{j}^{\circ})(1-b^{*}_{q|g}(z_{j}^{\circ}))]\geq[s^{\prime}_{g}(z_{i}^{\circ})b^{*}_{q|g}(z_{j}^{\circ})+s^{\prime}_{b}(z_{i}^{\circ})(1-b^{*}_{q|g}(z_{j}^{\circ}))]+[s^{\prime}_{g}(z_{j}^{\circ})b^{*}_{q|g}(z_{i}^{\circ})+s^{\prime}_{b}(z_{j}^{\circ})(1-b^{*}_{q|g}(z_{i}^{\circ}))]\iff\alpha\geq\beta+(b^{*}_{q|g}(z_{i}^{\circ})-b^{*}_{q|g}(z_{j}^{\circ}))(s^{\prime}_{b}(z_{i}^{\circ})-s^{\prime}_{b}(z_{j}^{\circ}))$. Observe, $(Z^{\prime}_{2,0})^{c}=[Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{g,b}\cup Z^{\prime}_{g,02}\cup Z^{\prime}_{02,b}\cup Z^{\prime}_{0,0}\cup Z^{\prime}_{2,2}]\cup[Z^{\prime}_{g,b}\cup Z^{\prime}_{0,b}\cup Z^{\prime}_{g,2}]$ Case 1 (L.4.1.1): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{g,b}\cup Z^{\prime}_{g,02}\cup Z^{\prime}_{02,b}\cup Z^{\prime}_{0,0}\cup Z^{\prime}_{2,2}$. Then $s^{\prime}_{b}(z_{i}^{\circ})-s^{\prime}_{b}(z_{j}^{\circ})=0\implies\alpha\geq\beta$. But this is a contradiction with the assumed $\alpha<\beta$, thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{g,g}\cup Z^{\prime}_{b,b}\cup Z^{\prime}_{g,b}\cup Z^{\prime}_{g,02}\cup Z^{\prime}_{02,b}\cup Z^{\prime}_{0,0}\cup Z^{\prime}_{2,2}$.
Case 2 (L.4.1.2): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,b}\cup Z^{\prime}_{0,b}\cup Z^{\prime}_{g,2}$. Then, $\alpha\geq\beta+(b^{*}_{q|g}(z_{i}^{\circ})-b^{*}_{q|g}(z_{j}^{\circ}))(s^{\prime}_{b}(z_{i}^{\circ})-s^{\prime}_{b}(z_{j}^{\circ}))\iff b^{*}_{q|g}(z_{i}^{\circ})\geq b^{*}_{q|g}(z_{j}^{\circ})\implies b^{*}_{q|g}(z_{i}^{\circ})(s^{\prime}_{g}(z_{i}^{\circ})-s^{\prime}_{g}(z_{j}^{\circ}))\geq b^{*}_{q|g}(z_{j}^{\circ})(s^{\prime}_{g}(z_{i}^{\circ})-s^{\prime}_{g}(z_{j}^{\circ}))\iff\alpha\geq\beta$ since $(s^{\prime}_{g}(z_{i}^{\circ})-s^{\prime}_{g}(z_{j}^{\circ}))\geq 0$ for $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{g,b}\cup Z^{\prime}_{0,b}\cup Z^{\prime}_{g,2}$. But this is a contradiction with $\alpha<\beta$, thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{g,b}\cup Z^{\prime}_{0,b}\cup Z^{\prime}_{g,2}$
Case 3 (L.4.1.3): $(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{2,0}$. Then, $(s^{\prime}_{b}(z_{i}^{\circ})-s^{\prime}_{b}(z_{j}^{\circ}))(b^{*}_{q|g}(z_{i}^{\circ})-b^{*}_{q|g}(z_{j}^{\circ}))=b^{*}_{q|g}(z_{i}^{\circ}))-b^{*}_{q|g}(z_{j}^{\circ}))\geq 0$ since $C_{2,0}(b^{*}_{q|g})=\emptyset$. Equivalently, $\alpha\geq\beta$, which is a contradiction with $\alpha<\beta$, thus $(z_{i}^{\circ},z_{j}^{\circ})\notin Z^{\prime}_{2,0}$. Thus, we have shown that $(z_{i}^{\circ},z_{j}^{\circ})\notin\mathcal{Z}\times\mathcal{Z}\implies b^{*}_{q|g}\notin A^{*}_{q|g}$.
Part 2: Next we show that $b^{*}_{q|g}\in A^{*}_{q|g}\implies b^{*}_{q|g}\in B^{*}_{q|g}$. $b^{*}_{q|g}\in A^{*}_{q|g}\implies\forall(z_{i},z_{j})\in\mathcal{Z}\times\mathcal{Z},\alpha=s^{\prime}_{g}(z_{i})b^{*}_{q|g}(z_{i})+s^{\prime}_{g}(z_{j})b^{*}_{q|g}(z_{j})\geq s^{\prime}_{g}(z_{i})b^{*}_{q|g}(z_{j})+s^{\prime}_{g}(z_{j})b^{*}_{q|g}(z_{i})=\beta$. Thus, $\forall(z_{i}^{\circ},z_{j}^{\circ})\in Z^{\prime}_{2,0}$, $\alpha=b^{*}_{q|g}(z_{i}^{\circ})\geq b^{*}_{q|g}(z_{j}^{\circ})=\beta\implies C_{2,0}(b^{*}_{q|g})=\emptyset\implies b^{*}_{q|g}\in B^{*}_{q|g}$.
Having shown both sides of the implication, we conclude $A^{*}_{q|g}=B^{*}_{q|g}$.
@@
(Maximal Sufficient Explainable Performance) $\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|g}$ where $A^{*}_{q|g}=\{a^{*}_{q|g}\colon a^{*}_{q|g}=\arg\max_{a^{*}_{q}\in A^{*}_{q}}\bar{t}^{\prime}_{g}(a^{*}_{q})\}$ and $\bar{t}^{\prime}_{g}(a^{*}_{q})=\frac{1}{n}\sum_{i=1}^{n}t^{\prime}_{g}(a^{*}_{q},z_{i})=\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a^{*}_{q}(z_{i})$
@@
From Lemma 4 it is sufficient to show that $a^{\prime}_{q}\in B^{*}_{q|g}$ or equivalently that $a^{\prime}_{q}\in A^{*}_{q}$ and that $\forall(z_{i},z_{j})\in Z^{\prime}_{2}\times Z^{\prime}_{0},a^{\prime}_{q}(z_{i})\geq a^{\prime}_{q}(z_{j})$.
From Proposition 1, $a^{\prime}_{q}\in A^{*}_{q}$.
From Lemma 1, $\tilde{r}^{\prime}(z_{i})>\tilde{r}^{\prime}(z_{j})\iff\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z_{i}))>\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z_{j}))\implies\mathbb{I}\{\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z_{i}))>n_{1-q}\}\geq\mathbb{I}\{\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z_{j}))>n_{1-q}\}\iff a^{\prime}_{q}(z_{i})\geq a^{\prime}_{q}(z_{j})$. Thus, $a^{\prime}_{q}\in A^{**}_{q}$. ∎
Proposition 3.
(Monotone Allocation) $\forall q_{i}<q_{j}\in\mathcal{Q}$, $\{z:z\in\mathcal{Z},a^{\prime}_{q_{i}}(z)=1\}\subseteq\{z:z\in\mathcal{Z},a^{\prime}_{q_{j}}(z)=1\}$
Proof.
From Lemma 2, $\{z:z\in\mathcal{Z},a^{\prime}_{q_{i}}(z)=1\}=\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z))>n_{1-q_{i}}\}$. Since $q_{i}<q_{j},n_{1-q_{i}}>n_{1-q_{j}}$, thus $\{z:z\in\mathcal{Z},a^{\prime}_{q_{i}}(z)=1\}=\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z))>n_{1-q_{i}}\}\subseteq\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(\tilde{r}^{\prime}(z))>n_{1-q_{j}}\}=\{z:z\in\mathcal{Z},a^{\prime}_{q_{j}}(z)=1\}$. ∎

### v3

@@ A. Theoretical Results
Assume for contradiction that $\exists b_{q}\in A^{*}_{q}$ s.t. $S(b_{q})>S(a^{\prime}_{q})$.
$\implies\exists v_{i},v_{j}\in V^{n}$ s.t. $s(b_{q},v_{i})+s(b_{q},v_{j})>s(a^{\prime}_{q},v_{i})+s(a^{\prime},v_{j}),b_{q}(v_{i})\neq a^{\prime}_{q}(v_{i}),b_{q}(v_{j})\neq a^{\prime}_{q}(v_{j})$
w.l.o.g. take $a^{\prime}_{q}(v_{i})>a^{\prime}_{q}(v_{j})$ and $b_{q}(v_{i})<b_{q}(v_{j})$
then $s(b_{q},v_{i})+s(b_{q},v_{j})=s_{b}(z_{i})+s_{g}(z_{j})$ and $s(a^{\prime}_{q},v_{i})+s(a^{\prime},v_{j})=s_{g}(z_{i})+s_{b}(z_{j})$
Now, taking it by case:
for $(z_{i},z_{j})\in Z_{k}\times Z_{k},k\in\{0,2,g,b\},s_{b}(z_{i})+s_{g}(z_{j})=s_{g}(z_{i})+s_{b}(z_{j})\Rightarrow\Leftarrow$
for $(z_{i},z_{j})\in Z_{g,0}\cup Z_{0,b}\cup Z_{g,b}\cup Z_{g,2}\cup Z_{2,b},s_{b}(z_{i})+s_{g}(z_{j})<s_{g}(z_{i})+s_{b}(z_{j})\Rightarrow\Leftarrow$
for $(z_{i},z_{j})\in Z_{0,g}\cup Z_{b,0}\cup Z_{b,g}\cup Z_{2,g}\cup Z_{b,2},r(z_{i})<r(z_{j})\implies a^{\prime}_{q}(z_{i})\leq a^{\prime}_{q}(z_{j})\Rightarrow\Leftarrow$
$\implies\nexists b_{q}\in A^{*}_{q}$ s.t. $S(b_{q})>S(a^{\prime}_{q})\implies a^{\prime}_{q}\in A^{*}_{q}$
@@
(Maximal Sufficient Explainable Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|g}$ where $A^{*}_{q|g}=\{a^{*}_{q|g}\colon a^{*}_{q|g}=\arg\max_{a_{q}\in A_{q}}S_{g}(a_{q})\}$
@@
Assume for contradiction that $\exists b_{q}\in A^{*}_{q|g}$ s.t. $S_{g}(b_{q})>S_{g}(a^{\prime}_{q})$
$\implies\exists v_{i},v_{j}\in V^{n}$ s.t. $s_{g}(z_{i})b_{q}(v_{i})+s_{g}(z_{j})b_{q}(v_{j})>s_{g}(z_{i})a^{\prime}_{q}(v_{i})+s_{g}(z_{j})a^{\prime}_{q}(v_{j}),b_{q}(v_{i})\neq a^{\prime}_{q}(v_{i}),b_{q}(v_{j})\neq a^{\prime}_{q}(v_{j})$
w.l.o.g. take $a^{\prime}_{q}(v_{i})>a^{\prime}_{q}(v_{j})$ and $b_{q}(v_{i})<b_{q}(v_{j})$
then $s_{g}(z_{i})b_{q}(v_{i})+s_{g}(z_{j})b_{q}(v_{j})=s_{g}(z_{j})$ and $s_{g}(z_{i})a^{\prime}_{q}(v_{i})+s_{g}(z_{j})a^{\prime}_{q}(v_{j})=s_{g}(z_{i})$
Now, taking it by case:
for $(z_{i},z_{j})\in Z_{k,k},k\in\{0,2,g,b\},s_{g}(z_{i})=s_{g}(z_{j})\Rightarrow\Leftarrow$
for $(z_{i},z_{j})\in Z_{b,0}\cup Z_{2,g}\cup Z_{0,b}\cup Z_{g,2},s_{g}(z_{i})=s_{g}(z_{j})\Rightarrow\Leftarrow$
for $(z_{i},z_{j})\in Z_{g,0}\cup Z_{2,0}\cup Z_{g,b}\cup Z_{2,b},s_{g}(z_{i})>s_{g}(z_{j})\Rightarrow\Leftarrow$
for $(z_{i},z_{j})\in Z_{0,g}\cup Z_{0,2}\cup Z_{b,g}\cup Z_{b,2},r(z_{i})<r(z_{j})\implies a^{\prime}_{q}(v_{i})\leq a^{\prime}_{q}(v_{j})\Rightarrow\Leftarrow$
$\implies\nexists b_{q}\in A^{*}_{q|g}$ s.t. $S_{g}(b_{q})>S_{g}(a^{\prime}_{q})\implies a^{\prime}_{q}\in A^{*}_{q|g}$
Proposition 3.
(Conditional Maximum Absolute Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q|s}$ where $A^{*}_{q|s}=\{a^{*}_{q|s}\colon a^{*}_{q|s}=\arg\min_{a_{q}\in A^{*}_{q}\cap A^{*}_{q|g}}L(a_{q})\}$,
$L(a_{q})=\frac{1}{n}\sum l(g(x_{i}),y_{i})a_{q}(v_{i})+l(b(x_{i}),y_{i})(1-a_{q}(v_{i}))$
@@
Assume for contradiction that $\exists b_{q}\in A^{*}_{q|s}$ s.t. $L(b_{q})<L(a^{\prime}_{q})$
$\implies\exists v_{i},v_{j}\in V^{n}$ s.t. $l_{i}(a^{\prime}_{q})+l_{j}(a^{\prime}_{q})>l_{i}(b_{q})+l_{j}(b_{q})$
Now, taking $(v_{i},v_{j})$ by case:
for $(v_{i},v_{j})\in Z_{m,n},m\neq n$, we have $a^{\prime}_{q}(v_{i})=b_{q}(v_{i}),a^{\prime}_{q}(v_{j})=b_{q}(v_{j})$, as both $a^{\prime}_{q},b_{q}\in A^{*}_{q}\cap A^{*}_{q|g}$
$\implies l_{i}(a^{\prime}_{q})+l_{j}(a^{\prime}_{q})=l_{i}(b_{q})+l_{j}(b_{q})\Rightarrow\Leftarrow$
for $(v_{i},v_{j})\in Z_{k,k}$, w.l.o.g. take $a^{\prime}_{q}(v_{i})>a^{\prime}_{q}(v_{j})$ and $b_{q}(v_{i})<b_{q}(v_{j})$
then $l_{i}(a^{\prime}_{q})+l_{j}(a^{\prime}_{q})>l_{i}(b_{q})+l_{j}(b_{q})\Longleftrightarrow l(g(x_{i}),y_{i})+l(b(x_{j}),y_{j})>l(b(x_{i}),y_{i})+l(g(x_{j}),y_{j})$
however, $a^{\prime}_{q}(v_{i})>a^{\prime}_{q}(v_{j})\implies r(z_{i})>r(z_{j})$
$\Longleftrightarrow-\sigma(l(g(x_{i}),y_{i})-l(b(x_{i}),y_{i}))>-\sigma(l(g(x_{j}),y_{j})-l(b(x_{j}),y_{j}))$
$\Longleftrightarrow l(g(x_{i}),y_{i})+l(b(x_{j}),y_{j})<l(b(x_{i}),y_{i})+l(g(x_{j}),y_{j})\Rightarrow\Leftarrow$
$\implies\nexists b_{q}\in A^{*}_{q|s}$ s.t. $L(b_{q})<L(a^{\prime}_{q})\implies a^{\prime}_{q}\in A^{*}_{q|s}$
@@
(Monotone Allocation)
$\forall q_{i}<q_{j}\in\mathcal{Q}$, $\{v:v\in\mathcal{V},a^{\prime}_{q_{i}}(v)=1\}\subseteq\{v:v\in\mathcal{V},a^{\prime}_{q_{j}}(v)=1\}$
@@
From Lemma 2, $\{v:v\in\mathcal{V},a^{\prime}_{q_{i}}(v)=1\}=\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(r(z))>n_{1-q_{i}}\}$. Then since $q_{i}<q_{j},n_{1-q_{i}}>n_{1-q_{j}}$, thus $\{z:z\in\mathcal{Z},a^{\prime}_{q_{i}}(z)=1\}=\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(r(z))>n_{1-q_{i}}\}\subseteq\{z:z\in\mathcal{Z},\mathrm{rank}_{D^{n}}(r(z))>n_{1-q_{j}}\}=\{v:v\in\mathcal{V},a^{\prime}_{q_{j}}(v)=1\}$. ∎
B. Dataset Details